If a balloon is filled with He, Ne, and Ar gases at a total pressure of 2.0atm, what is the partial pressure of argon if the partial pressures of helium and neon are both 0.25atm each?
1
Expert's answer
2020-04-03T11:52:22-0400

P=PHe + PNe + PAr


2=0.25+0.25+PAr

PAr=1.5


For more accurate calculations, use the formula

P =(nRT/V)He + (nRT/V)Ne + (nRT/V)Ar 


Please provide temperature and mass